# Symmetric quantum states: a review of recent progress

### v1

@@ Abstract
Symmetric quantum states are fascinating objects. They correspond to multipartite systems that remain invariant under particle permutations. This symmetry is reflected in their compact mathematical characterisation and also in their unique physical properties: they exhibit genuine multipartite entanglement and notable robustness against noise and perturbations. These features make such states particularly well-suited for a wide range of quantum information tasks. Here, we provide a pedagogic analysis of the mathematical structure and relevant physical properties of this class of states. Beyond the theoretical framework, it is essential to develop robust tools for certifying and verifying the properties of symmetric states in experimental settings. In this regard, we explore how standard techniques – such as quantum state tomography, Bell tests, and entanglement witnesses – can be specifically adapted for symmetric systems. Next, we provide an up-to-date overview of the most relevant applications in which these states outperform other classes of states in specific tasks. Specifically, we address their central role in quantum metrology, highlight their use in quantum error correction codes, and examine their contribution in computation and communication tasks. Finally, we present the current state of the art in their experimental generation, ranging from systems of cold atoms to implementations via quantum algorithms. We also review the most recent breakthroughs obtained in the different experimental realizations. Despite the significant progress made in recent years regarding the characterization and application of symmetric quantum states, several intriguing questions remain unsolved. We conclude this review discussing some of these open problems and outlining promising directions for future research.
@@ III.1.3 Absolute separability
We present here recent results concerning absolute separability (AS) within the symmetric subspace. AS denotes the set of separable quantum states that remain separable under the action of any global unitary transformation. Consequently, states of the AS set depend solely on their spectrum (i.e., eigenvalues) and, for this reason, this property is also referred to as “separability from spectrum". Progress in characterising the extremal points of the convex AS set has been quite limited. However, a renewed interest has recently emerged [26, 27, 28, 29], partly driven by the potential role of AS as a foundation for a restricted resource theory of entanglement. Since unitary transformations are the primary way of generating entanglement – whether through quantum circuits or quenches – AS states cannot be transformed into entangled one by unitary evolution. Sufficient conditions for absolute separability using the inverse of linear maps have recently been proposed [29] which involved only one or two eigenvalues, however, the complete characterization of the AS set remains an open problem. A related issue is the so-called absolute PPT (APPT) problem, which concerns the convex set of states that remain PPT under any unitary transformation. In contrast to the AS set, the APPT set is completely characterized in arbitrary dimensions using a hierarchy of linear matrix inequalities [30], though closed conditions cannot be obtained from them. Whether the AS and APPT sets coincide remains an open conjecture in entanglement theory.
@@ IV.1 Quantum state tomography
To conclude, PI QST enables reconstruction of PI states (and, in particular, of symmetric states) for larger number of qubits $N$. The advantages of this approach together with compressed sensing  [92] are acknowledged in [96] by analyzing a photonic 6-qubit Dicke state. Besides PI QST, additional techniques tailored to symmetric states include the use of optimal mutually unbiased bases [97]. Other recent advances in the field comprise classical shadows [98] and machine learning-assisted QST [99].
@@ IV.4 Entanglement certification
In the previous section we have seen that the cone of completely positive matrices, $\mathcal{CP}_{d}$, plays a crucial role for the separability of two-qudit DS states. Another interesting feature derives from considering its dual cone, known as the copositive cone, and denoted as $\mathcal{COP}_{d}$. Formally, a $d\times d$ matrix $H$ is said to be copositive if $\vec{x}^{T}H\vec{x}\geq 0$, for every entry-wise nonnegative vector $\vec{x}=(x_{1},\dots,x_{d-1})\in\mathbb{R}^{d}$. Similarly to the case of completely positive matrices, also copositive matrices find applications in a variety of problems, mostly related to the field of mathematical optimization (for a review of such problems see [125]). When it comes to entanglement detection, copositive matrices reveal an interesting connection to the concept of entanglement witness. An entanglement witness, $W\in\mathcal{H}$, is a Hermitian operator which satisfies two properties: i) $\mbox{Tr}(W\rho_{sep})\geq 0$ for every separable state $\rho_{sep}\in\mathcal{B}(\mathcal{H})$, and ii) there exists at least one entangled state $\sigma_{AB}\in\mathcal{B}(\mathcal{H})$ such that $\mbox{Tr}(W\sigma_{sep})<0$. From a geometrical point of view, as a consequence of the Hahn-Banach theorem, entanglement witnesses can be seen as hyperplanes separating the the set of the entangled states they detect from the convex set of separable states. In the context of entanglement detection, particularly relevant is the concept of decomposability for entanglement witnesses. If $W$ can be expressed as
@@
for some positive operators $P$ and $Q$, then $W$ is said to be decomposable; on the contrary, if such decomposition does not exist, it is dubbed non-decomposable. This latter class of entanglement witnesses is extremely interesting, as they are the sole candidates capable of detecting PPT-entangled states (for a comprehensive review of entanglement witnesses we refer the reader to [126]). Similarly to the case of entanglement witnesses, also copositive matrices are endowed with a natural concept of decomposition. Indeed, it is a known result that $\mathcal{PSD}_{d}+\mathcal{N}_{d}\subseteq\mathcal{COP}_{d}$, where $\mathcal{PSD}_{d}$ and $\mathcal{N}_{d}$ are the convex cones of positive semidefinite and of entry-wise nonnegative matrices of order $d$, respectively. In particular, it has been proven by Diananda that $\mathcal{COP}_{d}=\mathcal{PSD}_{d}+\mathcal{N}_{d}$ for $d\leq 4$, while the inclusion is strict for any $d\geq 5$ [127]. As a consequence, any copositive matrix $H\in\mathcal{COP}_{d}$ in $d\leq 4$ admits a decomposition of the form
@@ V.2.1 Quantum error correction
where $s,r\in{0,1}$, $A,B\in\{X,Y,Z\}$. Direct calculation shows how $M$ factorizes as per $M_{sA,rB}\propto\delta_{sr}\delta_{AB}$, where $\delta$ is the Kronecker symbol. The structure of the matrix $M$ tells us that it is possible to determine which type of error has occurred ($X$, $Y$ or $Z$) without destroying the coherence of the logical state and restore it by means of unitary transformations. The previous observation stems from the fact that the errors send $\{\left|0_{L}\right\rangle,\left|1_{L}\right\rangle\}$ to two-dimensional orthogonal subspaces corresponding to each type of error [168]. Checking how the system fixes elementary errors $\{X,Y,Z\}$ is sufficient for correctability of an arbitrary linear combination of them (i.e. a flip in a tilted Bloch orientation) or a statistical mixture of them (e.g. implemented by a Pauli channel). Moreover, by PI invariance, the code corrects equally well a flip in a different location $i\neq 1$, even if such location remains unknown.
Such class of codes has been extended to encode more than one qubit [169]. PI codes have been recently proposed for quantum deletion channels [170] and to correct any number of Pauli errors, deletions and spontaneous decay errors [171], which severely afflict photonic qubits and Rydberg atom experiments [172].
@@ VI.1.2 Atomic ensembles
So far, all these experiments have been carried out in spin-$1/2$ particles. Atomic ensembles of higher spin (particularly, spin-$1$, corresponding to three-level Dicke states) have also been experimentally investigated extensively [210, 211, 212, 213, 156]. Multilevel collective measurements in spinor BECs can be routinely implemented [214, 215] by splitting the population in each atomic level via a magnetic field and estimate the particle number via absorption imaging [211]. Measurements in other bases can be implemented via unitary transformations prior to measurement [216]. Such hyperfine spin states can be accessed via radio-frequency pulses and quadratic Zeeman energy shifts [156, 217]. Entanglement has been detected in such multilevel systems, whereas the experimental detection of Bell correlations in higher-dimensional many-body systems remains open.
@@ VI.5 Mechanical resonators in circuit quantum acoustodynamics
By means of phonon transmission, surface acoustic waves serve as quantum transducers that can propagate quantum information encoded in various quantum degrees of freedom [352, 353]. This phonon-mediated quantum state transfer allows for remote qubit entanglement [354] and for the creation and control of multi-phonon Fock states in bulk acoustic-wave resonators [355].
@@ VI.6 Quantum algorithms for Dicke state preparation
Dicke states can also be prepared in quantum computers via specifically designed quantum circuits of different complexity. There exist quantum algorithms that probabilistically prepare Dicke states in time $O(N\log N)$ with $O(\log N)$ ancillary qubits [360]. Deterministic algorithms for the preparation of qubit Dicke states $\left|D^{N}_{k}\right\rangle$ with complexity $O(kN)$ in the number of gates and $O(N)$ in quantum circuit depth, without ancillary qubits have been proposed in [361] such as the one that can be seen in Figure 8. These algorithms can prepare symmetric states in $O(N^{2})$ gates and $O(N)$ circuit depth. Follow-up works have investigated the preparation of Dicke states in cloud quantum computers [362], as well as divide-and-conquer approaches [363] and the effect of hardware connectivity limitations [364] ($O(k\log(N/k))$ for all-to-all connectivity vs. $O(\sqrt{kN})$ in a grid), which is an improvement over earlier proposals [365]. For specific classes, such as W states, the complexity can be reduced to $O(\log N)$ circuit depth and $O(N)$ entangling (CNOT) gates [366]. Similarly, dedicated quantum circuits for the preparation of qudit Dicke states have been recently proposed [367, 368, 369], which serve as preliminary steps for more general classes of quantum states related to integrable models, such as Bethe ansätze [370, 371, 372, 373, 374].
@@ VI.7 Dicke states in physical systems
as one tunes the external magnetic field $h$, in the isotropic phase $\gamma=1$, $\left|D^{N}_{k}\right\rangle$ appear as ground states of the LMG model for qubits. The LMG Hamiltonian is one of the few one-dimensional long-range quantum systems that has been exactly solved, including its entanglement entropy [390], and it was originally introduced in the context of nuclear physics, to describe a system of $n$ fermions in two levels, each $n$ times degenerate. Analogous studies have been carried out for generalized isotropic LMG models [391], where qudit Dicke states appear as its ground state for suitable parameter choices.
The LMG Hamiltonian can be understood as the effective Dicke Hamiltonian [1] when phonons have been adiabatically eliminated [275].
@@ VII Concluding remarks
Yet, despite this progress, several important open questions remain, pointing to promising avenues for future research. One such challenge is the general characterization of PPT-entangled symmetric states, both in bipartite and multipartite systems. In the case of bipartite diagonal symmetric (DS) states, it has recently been shown that separability is equivalent to a cyclic variation of the complete positivity problem [397], and PPT-entangled DS states have been fully characterized for dimension $d\leq 5$. For multipartite diagonal symmetric states much less is known: while for $N=3$ and $d\leq 4$ a complete solution has been provided in [22], already for $N\geq 4$, the complexity of this task becomes significantly harder, indicating the need for new mathematical tools to address this problem. Studying the properties of non-decomposable entanglement witnesses for symmetric states [398, 26] is also of interest. Aside from their connection with the cone of copositive matrices, advances in this area will led to the discovery of novel properties of positive, but not completely positive, maps. Symmetric states may also play a key role in addressing the conjecture, originally formulated in [399], that no PPT state in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ can have maximal Schmidt number (that is, equal to $d$). While the conjecture has been proven for $d=3$ [400], extending this result to higher dimension remains a challenging task. A closely related topic is the use of symmetric states to prove the PPT2 conjecture [401, 402]– stating that the composition of two PPT channels is entanglement breaking. For bipartite systems of qudits, this conjecture holds trivially for $d=2$, as a consequence of the PPT criterion. Moreover, in [403], it has been proven for $d=3$, but remains open for $d>3$. When restricting to symmetric states, it was shown in [404] that the conjecture holds true for Choi-type maps, which include bipartite DS states as a special case. However, an analogous result for symmetric, but not DS, states has yet to be established.

### v2

@@ Abstract
Symmetric quantum states are fascinating objects. They correspond to multipartite systems that remain invariant under particle permutations. This symmetry is reflected in their compact mathematical characterisation but also in their unique physical properties: they exhibit genuine multipartite entanglement and notable robustness against noise and perturbations. These features make such states particularly well-suited for a wide range of quantum information tasks. Here, we provide a pedagogic analysis of the mathematical structure and relevant physical properties of this class of states. Beyond the theoretical framework, robust tools for certifying and verifying the properties of symmetric states in experimental settings are essential. In this regard, we explore how standard techniques – such as quantum state tomography, Bell tests, and entanglement witnesses – can be specifically adapted for symmetric systems. Next, we provide an up-to-date overview of the most relevant applications in which these states outperform other classes of states in specific tasks. Specifically, we address their central role in quantum metrology, highlight their use in quantum error correction codes, and examine their contribution in computation and communication tasks. Finally, we present the current state-of-the-art in their experimental generation, ranging from systems of cold atoms to implementations via quantum algorithms. We also review the most significant results obtained in the different experimental realizations. Despite the notable progress made in recent years with regard to the characterisation and application of symmetric quantum states, several intriguing questions remain unsolved. We conclude this review by discussing some of these open problems and outlining promising directions for future research.
@@ III.1.3 Absolute separability
We present here recent results concerning absolute separability (AS) within the symmetric subspace. AS denotes the set of separable quantum states that remain separable under the action of any global unitary transformation. Consequently, states of the AS set depend solely on their spectrum (i.e., eigenvalues) and, for this reason, this property is also referred to as “separability from spectrum". Progress in characterising the extremal points of the convex AS set has been quite limited. However, a renewed interest has recently emerged [26, 27, 28, 29, 30], partly driven by the potential role of AS as a foundation for a restricted resource theory of entanglement. Since unitary transformations are the primary way of generating entanglement – whether through quantum circuits or quenches – AS states cannot be transformed into entangled one by unitary evolution. Sufficient conditions for absolute separability using the inverse of linear maps have recently been proposed [30] which involved only one or two eigenvalues, however, the complete characterization of the AS set remains an open problem. A related issue is the so-called absolute PPT (APPT) problem, which concerns the convex set of states that remain PPT under any unitary transformation. In contrast to the AS set, the APPT set is completely characterized in arbitrary dimensions using a hierarchy of linear matrix inequalities [31], though closed conditions cannot be obtained from them. Whether the AS and APPT sets coincide remains an open conjecture in entanglement theory.
@@ IV.1 Quantum state tomography
To conclude, PI QST enables reconstruction of PI states (and, in particular, of symmetric states) for larger number of qubits $N$. The advantages of this approach together with compressed sensing  [93] are acknowledged in [97] by analyzing a photonic 6-qubit Dicke state. Besides PI QST, additional techniques tailored to symmetric states include the use of optimal mutually unbiased bases [98]. Other recent advances in the field comprise classical shadows [99] and machine learning-assisted QST [100] as well as a novel approach, based on the symmetrization of observables, which allows for a significant reduction of the sample complexity [101].
@@ IV.4 Entanglement certification
The certification of entanglement, that is, deciding whether a given quantum state is entangled or not, can be realized in various ways. A typical approach relies on the construction of an entanglement witness (EW), that is, a Hermitian operator $W\in\mathcal{H}$ satisfying two properties: i) $\mbox{Tr}(W\rho_{sep})\geq 0$ for every separable state $\rho_{sep}\in\mathcal{B}(\mathcal{H})$, and ii) there exists at least one entangled state $\sigma\in\mathcal{B}(\mathcal{H})$ such that $\mbox{Tr}(W\sigma)<0$. From a geometrical point of view, as a consequence of the Hahn-Banach theorem, EWs can be seen as hyperplanes separating the the set of the entangled states they detect from the convex set of separable states. In the context of entanglement detection, particularly relevant is the concept of decomposability of an EW. If $W$ can be expressed as
@@
for some positive operators $P$ and $Q$, then $W$ is said to be decomposable; on the contrary, if such decomposition does not exist, it is dubbed non-decomposable. This latter class of entanglement witnesses is extremely interesting, as they are the sole candidates capable of detecting PPT-entangled states (for a comprehensive review of entanglement witnesses we refer the reader to [127]).
In the previous section, we have seen that the cone of completely positive matrices, $\mathcal{CP}_{d}$, plays a crucial role in the separability problem for two-qudit DS states. Another interesting feature derives from considering its dual cone, known as the copositive cone, and denoted as $\mathcal{COP}_{d}$. Formally, a $d\times d$ matrix $H$ is said to be copositive if $\vec{x}^{T}H\vec{x}\geq 0$, for every entry-wise nonnegative vector $\vec{x}=(x_{1},\dots,x_{d-1})\in\mathbb{R}^{d}$. Similarly to the case of completely positive matrices, also copositive matrices find applications in a variety of problems, mostly related to the field of mathematical optimization (for a review of such problems see [128]). Remarkably, copositive matrices reveal an interesting connection to the concept of entanglement witness. In fact, similarly to the case of entanglement witnesses, they are also endowed with a natural concept of decomposition. Indeed, it is a known result that $\mathcal{PSD}_{d}+\mathcal{N}_{d}\subseteq\mathcal{COP}_{d}$, where $\mathcal{PSD}_{d}$ and $\mathcal{N}_{d}$ are the convex cones of positive semidefinite and of entry-wise nonnegative matrices of order $d$, respectively. In particular, it has been proven by Diananda that $\mathcal{COP}_{d}=\mathcal{PSD}_{d}+\mathcal{N}_{d}$ for $d\leq 4$, while the inclusion is strict for any $d\geq 5$ [129]. As a consequence, any copositive matrix $H\in\mathcal{COP}_{d}$ in $d\leq 4$ admits a decomposition of the form
@@
In the case of many-body systems, EWs based on spin-squeezed inequalities can be used as a tool for entanglement certification. This approach is particularly suited to the case of many-body spin systems, for which the global spin operators can be defined as as $J_{\alpha}=\sum_{i=1}^{N}\frac{1}{2}\sigma_{i}^{\alpha}$, with $\alpha=\{x,y,z\}$. Hence, a state with total spin $J$ is said to be spin-squeezed [133] if
$\xi^{2}\equiv\frac{(\Delta J_{\bm{n}_{1}})^{2}}{\langle J_{\bm{n}_{2}}\rangle^{2}+\langle J_{\bm{n}_{3}}\rangle^{2}}<1~{},$ (28)
where $(\Delta J_{\bm{n}})^{2}\equiv\langle J_{\bm{n}}^{2}\rangle-\langle J_{\bm{n}}\rangle^{2}$ and $J_{\bm{n}}\equiv\bm{n}\cdot\bm{J}$, for a given direction $\bm{n}=(\bm{n}_{1},\bm{n}_{2},\bm{n}_{3})$. Remarkably, in [134], a criterion to detect entanglement in spin-squeezed states was found. Expanding on this result, the relation between spin-squeezing inequalities and entanglement in many-body systems was explored in many other works (see [135] and references therein). In [136, 137, 138], some inequalities were constructed for symmetric $N$-qubit states, providing necessary and sufficient criteria to certify two- and three-qubit genuine multipartite entanglement, while being only necessary for a higher number of qubits.
@@ V.2.1 Quantum error correction
where $s,r\in{0,1}$, $A,B\in\{X,Y,Z\}$. Direct calculation shows how $M$ factorizes as per $M_{sA,rB}\propto\delta_{sr}\delta_{AB}$, where $\delta$ is the Kronecker symbol. The structure of the matrix $M$ tells us that it is possible to determine which type of error has occurred ($X$, $Y$ or $Z$) without destroying the coherence of the logical state and restore it by means of unitary transformations. The previous observation stems from the fact that the errors send $\{\left|0_{L}\right\rangle,\left|1_{L}\right\rangle\}$ to two-dimensional orthogonal subspaces corresponding to each type of error [176]. Checking how the system fixes elementary errors $\{X,Y,Z\}$ is sufficient for correctability of an arbitrary linear combination of them (i.e. a flip in a tilted Bloch orientation) or a statistical mixture of them (e.g. implemented by a Pauli channel). Moreover, by PI invariance, the code corrects equally well a flip in a different location $i\neq 1$, even if such location remains unknown. More generally,$gnu$ codes can correct any number of Pauli errors as well as spontaneous decay and deletion errors [177]. Such class of codes has been extended to encode multiple qubit systems [178]. Recently, a generalization of several known PI codes has been proposed, where fewer qubits are needed  [179]. Finally, an approach based on certain polynomials has allowed to construct families of PI codes also for qudit systems [180].
@@ VI.1.2 Atomic ensembles
So far, all these experiments have been carried out in spin-$1/2$ particles. Atomic ensembles of higher spin (particularly, spin-$1$, corresponding to three-level Dicke states) have also been experimentally investigated extensively [218, 219, 220, 221, 164]. Multilevel collective measurements in spinor BECs can be routinely implemented [222, 223, 224] by splitting the population in each atomic level via a magnetic field and estimate the particle number via absorption imaging [219]. Measurements in other bases can be implemented via unitary transformations prior to measurement [225]. Such hyperfine spin states can be accessed via radio-frequency pulses and quadratic Zeeman energy shifts [164, 226]. Entanglement has been detected in such multilevel systems, whereas the experimental detection of Bell correlations in higher-dimensional many-body systems remains open.
@@ VI.5 Mechanical resonators in circuit quantum acoustodynamics
By means of phonon transmission, surface acoustic waves serve as quantum transducers that can propagate quantum information encoded in various quantum degrees of freedom [361, 362]. This phonon-mediated quantum state transfer allows for remote qubit entanglement [363] and for the creation and control of multi-phonon Fock states in bulk acoustic-wave resonators [364, 365].
@@ VI.6 Quantum algorithms for Dicke state preparation
Dicke states can also be prepared in quantum computers via specifically designed quantum circuits of different complexity. There exist quantum algorithms that probabilistically prepare Dicke states in time $O(N\log N)$ with $O(\log N)$ ancillary qubits [370]. Deterministic algorithms for the preparation of qubit Dicke states $\left|D^{N}_{k}\right\rangle$ with complexity $O(kN)$ in the number of gates and $O(N)$ in quantum circuit depth, without ancillary qubits have been proposed in [371] such as the one that can be seen in Figure 8. These algorithms can prepare symmetric states in $O(N^{2})$ gates and $O(N)$ circuit depth. Follow-up works have investigated the preparation of Dicke states in cloud quantum computers [372], as well as divide-and-conquer approaches [373] and the effect of hardware connectivity limitations [374] ($O(k\log(N/k))$ for all-to-all connectivity vs. $O(\sqrt{kN})$ in a grid), which is an improvement over earlier proposals [375]. Recently, a new deterministic quantum circuit for the generation of Dicke states was proposed, whose depth scales as ($O(\log(k)\log(N/k)+k)$ and which does not require the use of ancillary qubits [376]. For specific classes, such as W states, the complexity can be reduced to $O(\log N)$ circuit depth and $O(N)$ entangling (CNOT) gates [377]. Similarly, dedicated quantum circuits for the preparation of qudit Dicke states have been recently proposed [378, 379, 380], which serve as preliminary steps for more general classes of quantum states related to integrable models, such as Bethe ansätze [381, 382, 383, 384, 385, 386].
@@ VI.7 Dicke states in physical systems
as one tunes the external magnetic field $h$, in the isotropic phase $\gamma=1$, $\left|D^{N}_{k}\right\rangle$ appear as ground states of the LMG model for qubits. The LMG Hamiltonian is one of the few one-dimensional long-range quantum systems that has been exactly solved, including its entanglement entropy [402], and it was originally introduced in the context of nuclear physics, to describe a system of $n$ fermions in two levels, each $n$ times degenerate. Moreover, there has been considerable interest in studying several entanglement measures of the ground states of this model [403, 404]. Similar investigations have been carried out for generalized isotropic LMG models [405], where qudit Dicke states appear as its ground state for suitable parameter choices.
@@ VII Concluding remarks
Yet, despite this progress, several important open questions remain, pointing to promising avenues for future research. One such challenge is the general characterization of PPT-entangled symmetric states, both in bipartite and multipartite systems. In the case of bipartite diagonal symmetric (DS) states, it has recently been shown that separability is equivalent to a cyclic variation of the complete positivity problem [411], and PPT-entangled DS states have been fully characterized for dimension $d\leq 5$. For multipartite diagonal symmetric states much less is known: while for $N=3$ and $d\leq 4$ a complete solution has been provided in [22], already for $N\geq 4$, the complexity of this task becomes significantly harder, indicating the need for new mathematical tools to address this problem. Studying the properties of non-decomposable entanglement witnesses for symmetric states [412, 27] is also of interest. Aside from their connection with the cone of copositive matrices, advances in this area will led to the discovery of novel properties of positive, but not completely positive, maps. Symmetric states may also play a key role in addressing the conjecture, originally formulated in [413], that no PPT state in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ can have maximal Schmidt number (that is, equal to $d$). While the conjecture has been proven for $d=3$ [414], extending this result to higher dimension remains a challenging task. A closely related topic is the use of symmetric states to prove the PPT2 conjecture [415, 416]– stating that the composition of two PPT channels is entanglement breaking. For bipartite systems of qudits, this conjecture holds trivially for $d=2$, as a consequence of the PPT criterion. Moreover, in [417, 416], it has been proven for $d=3$, but remains open for $d>3$. When restricting to symmetric states, it was shown in [418] that the conjecture holds true for Choi-type maps, which include bipartite DS states as a special case. However, an analogous result for symmetric, but not DS, states has yet to be established.
@@
To conclude, we would like to acknowledge that, in spite of our best efforts to maintain an up-to-date list of references, certain works may have been inadvertently omitted, and for this, we offer our sincere apologies.